A pulsed laser of 10 ms is used for

welding. If energy delivered to the weld

is 1 joule, calculate the power of pulsed

laser in watt unit.


1
Expert's answer
2021-06-14T08:18:15-0400

"P=\frac Et=\frac{1}{10\cdot 10^{-3}}=100~W."